solve this problem through cramer's rule
y-x=1
3y+x=6

Answer by ewatrrr(24785)

Hi,
.
Solving this system of simultaneous equations using cramer's rule:
.
This system sets up well. The idea is to eliminate one of the variables by adding the two equations together. We Need to do no more, to do that in this ex.
y -x=1
3y+x=6
.
4y = 7
y = 7/4 or 1 3/4
.
From the first equation
y -1 = x
x = 3/4